OpenStudy (anonymous):

How would you solve 4x^2 – 6x – 10 = 0?

OpenStudy (anonymous):

You can factor it or use the quadratic formula.

OpenStudy (raden):

this can be factored

OpenStudy (anonymous):

How would the quadratic formula be used?

OpenStudy (anonymous):

\[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[ ax^2+bx+c \]

OpenStudy (anonymous):

Isnt 4x^2-6x-10 already in ax^2+bx+c form?

OpenStudy (anonymous):

Yes, so plug in the coeffecients into the quadratic formula and you'll get the solutions.

OpenStudy (anonymous):

so itd be x=-6+- sqrt 6^2-4*4810/2*4?

OpenStudy (anonymous):

You can't drop the negative signs.\[ a=4, b=-6, c=-10 \]

OpenStudy (anonymous):

okay so x=-6+-sqrt-6^2-4*4*-10/2*4?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

can you help me simplify that my calculator is being stupid...and isnt wanting to do sqrt or exponents.,../:

OpenStudy (anonymous):

Remember the order of operations. Do it one operation at a time.

OpenStudy (agent0smith):

^ you should probably use some parentheses when you write math like that. It's almost impossible to read. Note you can divide both sides of the equation by 2 first \[4x^2 – 6x – 10 = 0 \] after dividing it all by 2 becomes \[\large 2x^2 - 3x -5 = 0\] and this can be factored, you don't need the quadratic formula.

OpenStudy (anonymous):

@agent0smith I need it in quadratic formula thats what the teacher said

OpenStudy (anonymous):

Is -6+-sqrt36- -160/8 correct? @wio

OpenStudy (agent0smith):

Okay, then you just need to plug it all in. Just make sure you use order of operations correctly... I'm guessing by the way you typed it, that you did not. You can not enter it that way into a calculator. Use parentheses.

OpenStudy (anonymous):

where would i put parentheses?...ugh i hate math...

OpenStudy (agent0smith):

\[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a} \] \[x = \frac{\left( 6 \pm \sqrt{\left( 6^2 - 4*4*(-10) \right)} \right)} {2*4}\]

OpenStudy (anonymous):

so first you do (-6^2-4*4*(-10))? then multiply that by -6 then divide that by 2*4?

OpenStudy (agent0smith):

first you do (-6^2-4*4*(-10))? yes. then multiply that by -6 ... no, there's no multiplication by -6.

OpenStudy (anonymous):

then what would you do?

OpenStudy (agent0smith):

First simplify everything under the square root.

OpenStudy (anonymous):

-6^2=36 4*4*-10=-160 so sqrt36- -160?

OpenStudy (agent0smith):

Yes, now simplify the 36-(-160)

OpenStudy (anonymous):

Would you multiply that or add it since its a - -?

OpenStudy (agent0smith):

Add it

OpenStudy (agent0smith):

Two negatives make a positive.

OpenStudy (anonymous):

so 36+160=196?

OpenStudy (agent0smith):

Correct. So now it's \[x = \frac{\left( 6 \pm \sqrt{196} \right)} {2*4}\]

OpenStudy (anonymous):

Then would you multiply 2*4? and I still dont understand what to do with the -6+- thing.

OpenStudy (agent0smith):

The +- means plus OR minus. It really means: \[x = \frac{\left( 6 +\sqrt{196} \right)} {2*4}\] or \[x = \frac{\left( 6 - \sqrt{196} \right)} {2*4}\] But yes, multiply the 2 by 4.

OpenStudy (agent0smith):

Or take the square root of 196.

OpenStudy (anonymous):

so then x=(-6+sqrt196)/8.

OpenStudy (agent0smith):

Close, but you shouldn't have a negative 6 there, since it should be -b, which is -(-6).

OpenStudy (agent0smith):

and don't forget x=(6+sqrt196)/8 OR x=(6-sqrt196)/8.

OpenStudy (anonymous):

so itd turn into a positive 6? and the sqrt196 is 14,right?

OpenStudy (agent0smith):

Correct. Now you have x=(6+14)/8 or x=(6-14)/8

OpenStudy (anonymous):

do I solve both x=(6+14)/8 and x=(6-14)/8?

OpenStudy (agent0smith):

Correct.

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (agent0smith):

No prob :) did you get the two solutions for x?

OpenStudy (anonymous):

I got 2.5=(6+14)/8 and -1=(6-14)/8

OpenStudy (agent0smith):

Correct :)

OpenStudy (agent0smith):

So x = 2.5 and x=-1

OpenStudy (anonymous):

Yes I got that! :)) Thanks again